if a dice is thrown 5 times, what is the chance that an even number will come up exactly 3 times?

Answer by stanbon(75887)

if a dice is thrown 5 times, what is the chance that an even number will come up exactly 3 times?
-------
Binomial Problem with n = 5 and p(even) = 1/2
---
P(x = 3) = 5C3(1/2)^5 = 10/32 = 5/16
------------
Cheers,
Stan H.